OpenStudy (masonsurfs):

Can someone please help me solve a system using elimination? its for algebra 2.

OpenStudy (518nad):

yeah okay

OpenStudy (518nad):

give it ot me

OpenStudy (masonsurfs):

Okay one sec

OpenStudy (masonsurfs):

-2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5

OpenStudy (masonsurfs):

@518nad

OpenStudy (danjs):

you can add some multiples of any two equations together and have a variable drop out, say rewriting equation 3, as equation3+equation1

OpenStudy (masonsurfs):

I tried to do that but there were so many steps and I got confused @DanJS

OpenStudy (masonsurfs):

I've gotten a different answer every time

OpenStudy (danjs):

Here it is after doing eq3=eq3+eq1, and eq2=(-1)*eq2+eq1 -2x+2y+3z=0 3y+2z=3 5y+6z=5 ------------------

OpenStudy (masonsurfs):

So those are the same equations, just multiplied by -1?

OpenStudy (danjs):

yeah, i multiplied equation 2 by -1, then added equation 1 to it -2x+2y+3z=0 3y+2z=3 5y+6z=5 ------------------

OpenStudy (masonsurfs):

Sorry I'm still kinda lost... the first equation you wrote is equation one, correct?

OpenStudy (masonsurfs):

And how did you get the next two equations?

OpenStudy (danjs):

ok, when you have a system of linear equations like that, you can do 3 things. 1 - change the order of the equations 2 - multiply an equation by a constant number 3 - add equations together

OpenStudy (masonsurfs):

Okay

OpenStudy (masonsurfs):

So which one is the best option?

OpenStudy (danjs):

So starting from the given, you want to add equations together so that one of the variables goes to zero and drops out. -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 -------------- Rewrite equation3, as equation3+equation 1, so that the -2x+2x, will go to 0x and x will dissapear, leaving an equation in only y and z. -2x+2y+3z=0 -2x-y+z=-3 5y+6z=5 --------------- good on that so far?

OpenStudy (masonsurfs):

I understand the first part, but how did you get 5y+6z=5 ?

OpenStudy (danjs):

that is the new 3rd equation. it is now eq1+eq3, adding equation 1 to equation 3 gives that

OpenStudy (masonsurfs):

Okay I understand that part now

OpenStudy (danjs):

now do the same sorta thing to eliminate the x variable from either eq1 or eq 2

OpenStudy (masonsurfs):

Okay one sec

OpenStudy (masonsurfs):

So if I multiply eq1 by -1 then add that to eq2, it will cancel out the x value?

OpenStudy (danjs):

-2x+2y+3z=0 -2x-y+z=-3 5y+6z=5 --------------- you see you need to multiply one of them by -1 first, so the x terms add up to zero -1 times equation 2 gives -2x+2y+3z=0 2x+y-z=3 5y+6z=5 --------------- and then do equation 2 goes to the sum equation 2 plus equation 1, Eq2--->Eq2 + Eq1 -2x+2y+3z=0 3y+2z=3 5y+6z=5 ---------------

OpenStudy (danjs):

yeah you got it

OpenStudy (masonsurfs):

Okay I'm just reading it all one sec

OpenStudy (danjs):

You just have to remember what you can do, those 3 rules, and try then to use those to eliminate variables from equations.

OpenStudy (masonsurfs):

Okay thanks for all the help

OpenStudy (danjs):

You can do the same sort of thing to get rid of y or z in eq2 or eq3. This will leave an equation with 1 variable.

OpenStudy (danjs):

-3 times equation 2 -2x+2y+3z=0 -9y-6z=-9 5y+6z=5 --------------- Eq3 becomes eq3+eq2 -2x+2y+3z=0 3y+2z=3 -4y=-4 ---------------

OpenStudy (danjs):

Now you have enough to solve for each variable, the third equation solves to y=1, then use that to get the x and z values

OpenStudy (masonsurfs):

Okay I will try

OpenStudy (masonsurfs):

So would y=1? @DanJS

OpenStudy (danjs):

yes, from equation 3 , -4y=-4, y=1

OpenStudy (danjs):

use y=1 in equation 2 now, to see that z will be zero 3y+2z=3 ----- 3*1 + 2z = 3 z=0

OpenStudy (danjs):

now use those two y=1 and z=0, in equation 1 to get x -2x+2y+3z=0 -2x+2*1+3*0=0 -2x=-2 x=1 x also is 1, so the three variables are x=1 y=1 z=o you can test by putting those values into each equation, all 3 equations should hold true

OpenStudy (masonsurfs):

Okay, I understand now. Thank you so much it was very helpful. @DanJS

OpenStudy (danjs):

welcome